OpenStudy (cp9454):

A body is thrown with velocity 100m/s, it travels 5m in last second of its upward journey. if the same body is thrown with velocity 200m/s, what distance will it travel in last second of its upward journey?

OpenStudy (cp9454):

A. 5 B. 10 C. 20 D. 25

OpenStudy (anonymous):

To me, it 's A

OpenStudy (cp9454):

how, please explain.

OpenStudy (anonymous):

nvm, I really don't know how to explain @tkhunny

OpenStudy (tkhunny):

Location: \(s_{0}(t) = -4.9t^{2} + V_{0}t + H_{0}\) Location 100: \(s_{1}(t) = -4.9t^{2} + 100t + H_{1}\) Location 200: \(s_{2}(t) = -4.9t^{2} + 200t + H_{2}\) When does the first one reach its peak? \(t_{Peak1} = -b/(2a) = -100 / -9 = 11.1111111\) \(s_{1}(11.111) = -4.9(11.111)^{2} + 100(11.111) + H_{1}\) 1 Second sooner \(s_{1}(10.111) = -4.9(10.111)^{2} + 100(10.111) + H_{1}\) This will help you find \(H_{1} = H_{2}\) Location 200: \(s_{2}(t) = -4.9t^{2} + 200t + H_{2}\) Again, the maximum height is: \(t_{Peak2} = -b/(2a) = -200 / -9 = 22.2222222\) Maximum height is: \(s_{2}(22.222) = -4.9(22.222)^{2} + 200(22.222) + H_{2}\) One second sooner \(s_{2}(21.222) = -4.9(21.222)^{2} + 200(21.222) + H_{2}\) That's about it. Just walk your way through it, one piece at a time.

OpenStudy (tkhunny):

I was bored. Do you see my errors? First, -b/(2a) = 100/9.8, not 100/9. In any case, at some point, the faster one slows to 100 m/s and then follows the same trajectory as the slower one. Thus, they are the same at the peak. 5 is 5.

OpenStudy (anonymous):

Yes!! I am calculating and figure out the mistake there. hihihi Thanks for neat explanation. I got it.